OpenStudy (anonymous):

Solve for x: (x - 9)^2 = 1

OpenStudy (saifoo.khan):

8,10

OpenStudy (anonymous):

Both?

OpenStudy (saifoo.khan):

Yes.

OpenStudy (anonymous):

Oh okay thank you :D

OpenStudy (saifoo.khan):

Np. xD

OpenStudy (anonymous):

x-9=1 and -1

OpenStudy (jamesj):

The reason is that if y^2 = 1, then y = +1 or -1 Hence x-9 = 1 or x - 9 = -1 x = 10 x = 8

OpenStudy (anonymous):

Oh okay that makes sense. Thanks, JamesJ!